Find the first five terms of the recursively defined infinite sequence.
a1=2, ak+1=3ak-4
a1=
a2=
a3=
a4=
a5=

Answer by stanbon(75887)

Find the first five terms of the recursively defined infinite sequence.
a1=2
a(k+1) = 3a(k)-4
----------------------
a1 = 2
a2=3*2-4 = 2
a3=2
a4=2
a5=2
===========
Cheers,
Stan H.
===========